by boat, joseph can go 72km upstream to a fishing hole in 4 hours. to go back, he needs only 3 hours. what is his rate upstream and his rate downstream?: by boat, joseph can go 72km upstream to a fishing hole in 4 hours. to go back, he needs only 3 hours. what is his rate upstream and his rate downstream?
Answer by edjones(2120)

rt=d
r=d/t
72/4=18 km/h rate upstream
72/3=24 km/h rate downstream
.
Ed